In 1993, the life expectancy of males in a certain country was 61.9 years. In 2000, it was 65.6 years. Let E represent the life expectancy in year t and let t represent the number of years since 1994.

E(t)=_t+_ (round to the nearest tenth
Use the function to predict the life expectancy in 2005
E(12)=_

Answer by stanbon(75887)

In 1993, the life expectancy of males in a certain country was 61.9 years.
Let x = 0 represent 1993.
-----
In 2000, it was 65.6 years. Let E represent the life expectancy in year t and let t represent the number of years since 1994.
------
You have 2 points: (0,61.9), (7,65.6)
----
slope = (65.6-61.9)/(7-0) = 0.5286
-----
E-intercept is E(0) = 61.9
-----
E(t)= 0.53t+61.9_ (round to the nearest tenth
--------------------
Use the function to predict the life expectancy in 2005
E(12)= 0.53*12+61.9 = 68.26
---------------------------------
Cheers,
Stan H.
=========